Prev Next

SW&IT Meeting Minutes 2021-10-18

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20211018
LR20211004 - Bec, Matthieu 80.0
LR20211004 - Mechtley, Brandon 72.0
LR20211004 - Eychaner, Glenn 40.0
LR20211004 - Edmeades, David 80.0
LR20211004 - Grenz, Paul 20.0
LR20211004 - Hooper, Stephen 80.0
LR20211004 - Zhang, Yang 80.0
LR20211004 - Zhang, Xianyu 73.0

News from management

  • Performance Appraisal 2021
New forms and schedule https://hr.arizona.edu/career-conversations

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • OC, Management
    • AO anemometer trouble shooting
    • OCAM2k lab tests discussion with ENG and IT
    • learning some more about AO s/w, join discussions and scheming on #ao-sw
    • discuss /repository ro, permission, retention, register.cl script
  • This week
    • git/github "policy" (= philosophy slides) for new staff
    • review TCS punch-list from JHill with PGrenz
    • continued AO/dms discussions with Al
    • close swg recruitment
    • IT expenditures go-ahead
    • Performance Appraisals

DE

  • MS Server, again. Finally got a reply from Ruben this morning, asking for a copy of the reciept. (I thought he'd have it?) but that's progress.
  • Backup server is a sticky wicket, keeps getting into a bad state with updates. Have some more things to try
  • JP conversation with Google. She mentioned talks with a transition contractor. She's keeping me updated
  • MS Solidworks
  • New Zoom room for LF/SHARK, will hear from her soon but should be pretty standard.
  • Mountain repository cleaning script seems to be working again, AC is leading the conversation about trimming strategies that I brought up. No word from archive team re: error handling/notification
  • Quotes for new storage equipment

YZ

  • Previous:
    • Continue MODS WCS program (some revision, github and RPM setup, verification with Olga).
    • lbc-alerts list and LBC alert scripts.
    • TMS meeting and software work.
    • TMS data transfer utility (finished program, to be setup on robs).
    • Safety training.
  • Next:
    • Setup TMS data transfer utility on robs.
    • lbc-alerts list revision.
    • Hopefully finish MODS WCS verification.
    • TMS meeting and routine work.
    • Other misc LBC/MODS stuff.
    • Possible TCS and Argos swingarm test/debug with Paul.

GE

  • Last week
    • Digging through the 550,000 lines of code of the LUCI system.
    • Took ownership of IT 8487 and 8494, noting that these are more properly Aladin issues rather than RTD issues.
    • Met regarding LUCI image cubes, sub-frames, and AO requirements.
    • Read GEIRS manual and looked into using GEIRS to save a summed image along with each cube. This would prevent large cubes from freezing Aladin.
    • Dug through the LUCI configuration system after discovering that some parts of the config in my copied VM were hard-coded to isola-luci.tucson.lbto.org.
      • You can really mess up the LUCI configuration by logging in as an engineer and editing it. And finding one particular element of the config is not easy.
  • This week
    • Digging through the 550,000 lines of code of the LUCI system.
    • Prepare a new LUCI version for testing and deployment that will hopefully fix many of the open RTD issues (which are really Aladin issues).
    • Meet regarding LUCI sub-frame and AO requirements.

BM

  • Last week
    • Testing pause/resume changes to WfsArbitrator / AdSecArbitrator
    • Updated savestatus command in AdSecArbitrator to not read or save accelerometer data.
    • SX anemometer debugging
    • Listened in on a little bit of Keck AO planning workshop.
  • This week
    • Deployment / testing of new pause/resume commands on LBTI.
    • Logging fixes
      • Logging telemetry during SBCU timeout.
      • Make exception handling during connection in AnemometerMon less vague and make sure tcp connection gets closed on config failure.
      • Telemetry logging, fix field width
    • AO data migration

SH

  • SOUL computer
  • Anemometer
  • base camp camera with LeRoy
  • communications workshop on Thursday and Friday

next:

  • JH, linuxapps
  • communications workshop
  • quotes for IT hardware

PSG

  • Last week
    • Research connection limit on HL2000 device.
    • Supports up to 4 connections, can be set to any value between 1 and 4, inclusive.
  • This week
    • Cannot test this week due to the AO system being worked on.
    • Looking at TCS code to start to undrstand co-pointing errors.
    • Schedule time to test next week with Yang and Gustavo.
    • Taking vacation on Thursday and Friday (also Monday next week)
Topic revision: r9 - 19 Oct 2021, PaulGrenz
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback